Find y if the line through A(-1, 2) and B(3, 5) is parallel to the line through (0, -1) and (6, y).
Answer by richwmiller(17219)

(y1-y2)/(x1-x2)=m1 slope
y1=2
y2=5
x1=-1
x2=3
plug in those values and solve for m
Since the lines are parallel they have the same slope
do the same for line two
(y1-y2)/(x1-x2)=m2 slope
but we now know m1 and m2 but don't know y2
x1=0
x2=6
y1=-1
y2=y
another way is to set the two equal to each since m1=m2
(2-5)/(-1-3)=(-1-y)/0-6
-3/-4=-1-y/-6
3/4=(-1-y)/-6
mulitply both sides by -6
-18/4=-1-y
-9/2=-1-y
-4.5=-1-y
add 1 to both sides
-3.5=-y
add y to both sides and add 3.5 to both sides
OR multiply both sides by -1
y=3.5
check
-1-3.5/-6=-4.5/-6=3/4
ok